# Dual Language Models: Balancing Training Efficiency and Overfitting Resilience

### v1

@@ 7 Conclusion
In this work, we addressed the fundamental trade-off between the training efficiency of autoregressive models and the overfitting resilience of masked-diffusion models. We have empirically demonstrated that a dual-objective training strategy successfully achieves the best of both worlds, resulting in models that converge rapidly without any performance degradation in data-constrained settings. We established that combining objectives is universally beneficial and derived practical guidelines for selecting the optimal training ratio based on the degree of data repetition. We showed that prefix language modeling is induced and that it performs better than autoregressive evaluation on downstream tasks. While training on hundreds of data repetitions may seem extreme today, the asymmetry between exponentially scaling compute budgets and the finite supply of high-quality text suggests that data constraints will become increasingly relevant for frontier model development. Our findings indicate that dual-objective training provides a robust and compute-efficient path forward as the field approaches these fundamental limits.
@@
To ensure reproducibility of our work we provided the guidelines on how to train language models on both objectives at the same time in Section 3. For our model parameters and hyperparameters we specified those in Section 5.1. We describe how we perform the evaluations, the number of mask tokens used for PLL, the prompt formats, and log-likelihood normalizations in Section 4, Appendix B, and Appendix D. We openly release our custom training and evaluation code at https://github.com/ltgoslo/dual-language-models. The training code is based on the common and freely distributed deep-learning framework PyTorch (10.5555/3454287.3455008).
@@ Acknowledgments
This work is fully funded by the University of Oslo. The computations were performed on resources provided through Sigma2 – the national research infrastructure provider for high-performance computing and large-scale data storage in Norway. We acknowledge Norway and Sigma2 for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through project 465001890.
@@ Appendix A The use of large language models
Appendix B Log-likelihood normalization
For the BLiMP task, which is not considered in the OLMES evaluation suite, we do not apply any normalization and take the raw log-likelihood. We also stick to the no-context form of this task, where the whole sentence is considered the completion. We apply character length normalization to ARC-Easy, HellaSwag, MMLU, PIQA, and SIQA. Finally, we apply point-wise mutual information normalization (holtzman-etal-2021-surface), where the log-likelihood of the context-informed completion is divided by the log-likelihood of the uncontrained context completion, this can be seen in Equation 5, to ARC-Challenge, Commonsense QA, and OpenBook QA.
@@ Appendix B Log-likelihood normalization
Appendix C Monte Carlo estimation of log-likelihood
To evaluate the masked-diffusion capabilities of our models, we use Equation 4 with the same modification as for the autoregressive evaluation as well as an adaptation of Monte-Carlo sampling to estimate the log-likelihood of each completion. Instead of taking the expectation over $t\sim\mathcal{U}(0,1)$, we take the expectation between $N$ equally spaced points between 0 and 1. This reduces the variance of the estimation and allows for a faster convergence. However, accurate estimation still requires $N\geq 256$, which is unbearably slow – especially when compared to simple autoregressive calculation of log-likelihood that requires only a single forward pass.
Appendix D Pseudo log-likelihood estimation
@@
$\log p_{\bm{\theta}}(\bm{w})=\sum_{i=1}^{|{w}|}\log p_{\bm{\theta}}\bigl(w_{i}\mid\bm{c}$ $\oplus w_{0}\oplus\cdots\oplus w_{i-1}$ (6)
$\oplus\mathord{\mathrm{[MASK]}}$
@@
$\log p_{\bm{\theta}}(\bm{w})=\sum_{i=1}^{|\bm{w}|}\log p_{\bm{\theta}}\bigl(w_{i}\mid\bm{c}$ $\oplus w_{0}\oplus\cdots\oplus w_{i-1}$ (7)
$\oplus\operatorname{[MASK]}\oplus\cdots\oplus\operatorname{[MASK]}$
@@
where $n$ represents the number of $\operatorname{[MASK]}$ tokens. In our case, we take a combination of two different numbers of mask tokens (1 and 6), by taking the best score of the two for each task. The two values were chosen experimentally, more details on the results of each number of mask tokens can be found in Appendix G.
Appendix E Proof of left-shift closure
@@
The Restricted Access Sequence Processing language (RASP; pmlr-v139-weiss21a) is a sequence processing language that uses two types of variables: sequence operators and selectors; and two types of operators: element-wise and select-aggregate operators. Valid programs in RASP are operations on sequence operators formed by a finite composition of element-wise and select-aggregate operators.
@@
For every valid program written in RASP, there exists an equivalent fully-bidirectional transformer model that computes the same per-position operation; see pmlr-v139-weiss21a; 10.5555/3666122.3667771.
@@
Theorem 1 implies that when we parameterize a masked-diffusion model with a shifted transformer, it is as expressive as the standard non-shifted parameterization. More specifically, masked diffusion is defined in Equation 4, and $p_{\bm{\theta}}(x_{i}\mid\bm{x}^{t})$ is typically implemented as a fully-bidirectional transformer model that outputs this probability at the $i$th position. When we set $\Sigma$ as our subword vocabulary, we get that the space of all possible transformer realizations of $p_{\bm{\theta}}(x_{i}\mid\bm{x}^{t})$ are the $\Sigma$-realizable functions $\mathcal{R}_{\Sigma}$ (Definition 2). Theorem 1 shows that if we instead expect the output at the (i1)(i-1)th position, we do not lose any expressivity. Thus, transformer-based dual language models are a generalization of standard masked-diffusion language models. Note that the left-shift closure in Theorem 1 works up to the first token – which is guaranteed to be the special <s> token in Definition 2 as well as in the actual implementation.
Appendix F Validation loss curves
@@
Figure 6: Validation loss curves for 128 repetitions. These plots clearly demonstrate how training runs with high autoregressive ratio (in red) overfit. High masked-diffusion ratios are in blue.
@@
Figure 7: Validation loss curves for 4 repetitions. All losses monotonically decrease because overfitting is not a concern in this setting. High autoregressive ratios are plotted in red and high masked-diffusion ratios are shown in blue.
Appendix G Effects of number of mask tokens on the PLL
We first look at whether using a single number of mask tokens can lead to a good estimation of the PLL in general. For this, we evaluate five different models from 1 to 6 mask tokens and report the results in Tables 3, 4, 5, 6 and 7.
Table 3: PLL performance depending on the number of mask tokens. We show the PLL performance on the 9 tasks of the model trained with an equal ratio of masked-diffusion and AR and 32 repetitions with different number of masks. Best results per task are boldfaced.
@@ Appendix G Effects of number of mask tokens on the PLL
Table 4: PLL performance depending on the number of mask tokens. We show the PLL performance on the 9 tasks of the model trained with a 1 masked-diffusion to 7 autoregressive ratio and 32 repetitions with different number of masks. Best results per task are boldfaced.
@@
Table 5: PLL performance depending on the number of mask tokens. We show the PLL performance on the 9 tasks of the model trained with a 7 masked-diffusion to 1 autoregressive ratio and 32 repetitions with different number of masks. Best results per task are boldfaced.
@@
Table 6: PLL performance depending on the number of mask tokens. We show the PLL performance on the 9 tasks of the model trained with an equal ratio of masked-diffusion and AR and 16 repetitions with different number of masks. Best results per task are boldfaced.
@@
Table 7: PLL performance depending on the number of mask tokens. We show the PLL performance on the 9 tasks of the model trained with an equal ratio of masked-diffusion and AR and 64 repetitions with different number of masks. Best results per task are boldfaced.
@@
Table 8: PLL performance for combinations of one mask token and multi-mask token. Best results per model are bolfaced.
@@
Appendix H PLL versus Masked-Diffusion
Table 9: Normalized PLL versus Masked-Diffusion evaluation. The scores for each task are normalized so that 0% corresponds to the random baseline and 100% is the perfect score. The best result for each task is in boldfaced. We evaluate a model trained with equal AR and masked-diffusion ratio and 32 repetitions.
@@ Appendix H PLL versus Masked-Diffusion
Appendix I Prefix versus Autoregressive on optimal models.
@@
Table 10 shows that evaluting with the prefix mask almost always outperforms using the causal mask when the models are optimally trained. This is true in both the regular and constrained data settings.
Appendix J Detailed results of diffusion-masked evaluation
Table 11: The normalized PLL performance of selected models. We show the results on all nine evaluated tasks for three repetition values; each repetition group contains the results of the best-performing autoregressive-diffusion ratio and of the autoregressive-only model. The scores for each task are normalized so that 0% corresponds to random baseline and 100% is the perfect score. The best result for each dataset size is boldfaced.
@@
   Dual ($3:1$) 06.0 28.3 62.7 33.4 27.8 04.3 12.3 37.4 15.4 25.3
   Masked-Diffusion ($0:1$)  -0.1 22.3 64.8 29.0 24.1 01.6 09.1 27.2 14.4 21.4
@@
   Dual ($1:7$) 02.8 23.3 63.5 30.5 25.0 02.1 12.8 31.8 15.2 23.0
@@
Table 11 shows similar trends to those found in Table 2. The notable exception being for BLiMP where the performances are similar between both models. Unlike the autoregressive models, the performance of the purely masked-diffusion models is similar to each other. This is partially due to the model not overfitting, but also to it not being sample efficient. On the other hand we see that for the Dual Models, the performance significantly increases as we increase the training data set size.

### v2

@@ 7 Conclusion
In this work, we addressed the fundamental trade-off between the training efficiency of autoregressive models and the overfitting resilience of masked-diffusion models. We have empirically demonstrated that a dual-objective training strategy successfully achieves the best of both worlds, resulting in models that converge rapidly without any performance degradation in data-constrained settings. Crucially, because this unification requires no architectural changes, the resulting models incur no inference overhead and can be deployed as standard autoregressive transformers.
We established that combining objectives is universally beneficial and derived practical guidelines for selecting the optimal $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} based on the degree of data repetition. Furthermore, we observed that the diffusion objective induces robust prefix language modeling capabilities, leading to superior performance on downstream tasks compared to standard autoregressive baselines. While training on hundreds of data repetitions may seem extreme today, the asymmetry between exponentially scaling compute budgets and the finite supply of high-quality text suggests that data constraints will become increasingly relevant for frontier model development. Our findings indicate that dual-objective training provides a robust and compute-efficient path forward that retains standard inference capabilities as the field approaches these fundamental limits.
@@
To ensure reproducibility of our work we provided the guidelines on how to train language models on both objectives at the same time in Section 3. For our model parameters and hyperparameters we specified those in Section 5.1. We describe how we perform the evaluations, the number of mask tokens used for PLL, the prompt formats, and log-likelihood normalizations in Section 4, Appendix C, and Appendix E. We openly release our custom training and evaluation code at https://github.com/ltgoslo/dual-language-models. The training code is based on the common and freely distributed deep-learning framework PyTorch (Paszke et al., 2019). The trained models are released openly under the Apache 2.0 license for further investigation at https://huggingface.co/ltg/dual-lm-470m.
@@ Acknowledgments
The computations were performed on resources provided through Sigma2 – the national research infrastructure provider for high-performance computing and large-scale data storage in Norway. We acknowledge Norway and Sigma2 for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through project 465001890.
The efforts described in this paper were jointly funded by the University of Oslo and the HPLT project (High Performance Language Technologies; coordinated by Charles University).
@@ Appendix A The use of large language models
Appendix B Erratum: Loss formulation
The original version of this paper as well as the training code used the following formulation the loss function:
$\mathop{\text{argmin}}_{\bm{\theta}}\mathop{\mathbb{E}}_{\raisebox{-0.3pt}{${}_{\bm{x}\,\sim\,\mathcal{D}}$}}\Bigl[{2\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9}\mathcal{L}_{\text{AR}}(}\bm{x}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9};}\,\bm{\theta}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9})}+{\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}(1-\alpha)}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9}\mathcal{L}_{\text{MD}}(}\bm{x}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9};}\,\bm{\theta}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9})}\Bigl].$ (6)
The additional factor of two for $\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9}\mathcal{L}_{\text{AR}}$} was supposed to balance the $\nicefrac{{1}}{{t}}$ term applied when computing $\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9}\mathcal{L}_{\text{MD}}$} (recall Equation 4). This is however not mathematically correct (Equation 4) and unnecessary; the loss function is thus simplified in this updated version.
Appendix C Log-likelihood normalization
For the BLiMP task, which is not considered in the OLMES evaluation suite, we do not apply any normalization and take the raw log-likelihood. We also stick to the no-context form of this task, where the whole sentence is considered the completion. We apply character length normalization to ARC-Easy, HellaSwag, MMLU, PIQA, and SIQA. Finally, we apply point-wise mutual information normalization (Holtzman et al., 2021), where the log-likelihood of the context-informed completion is divided by the log-likelihood of the uncontrained context completion, this can be seen in Equation 7, to ARC-Challenge, Commonsense QA, and OpenBook QA.
@@ Appendix C Log-likelihood normalization
Appendix D Monte Carlo estimation of log-likelihood
To evaluate the masked-diffusion capabilities of our models, we use Equation 4 with the same modification as for the autoregressive evaluation as well as an adaptation of Monte-Carlo sampling to estimate the log-likelihood of each completion. Specifically, we estimate the following expected value:
$\int_{0}^{1}\mathop{\mathbb{E}}_{\bm{x}^{t}\sim q_{t\mid 0}(\cdot\mid\bm{x})}\Biggl[\frac{1}{t}\sum_{\left\{i\mid x_{i}^{t}=\,\texttt{mask}\right\}}\!\!\!\!\!\!\!\!\!\log p_{\bm{\theta}}(x_{i}\mid\bm{x}^{t})\Biggr]\,\mathrm{d}t=\mathop{\mathbb{E}}_{\begin{subarray}{c}t\sim\mathcal{U}(0,1)\\ \bm{x}^{t}\sim q_{t\mid 0}(\cdot\mid\bm{x})\end{subarray}}\Biggl[\frac{1}{t}\sum_{\left\{i\mid x_{i}^{t}=\,\texttt{mask}\right\}}\!\!\!\!\!\!\!\!\!\log p_{\bm{\theta}}(x_{i}\mid\bm{x}^{t})\Biggr]$ (8)
To reduce the variance of the estimation and get faster convergence, we take the expectation between $N$ equally spaced points between 0 and 1. instead of taking the expectation over $t\sim\mathcal{U}(0,1)$. Yet, accurate estimation still requires $N\geq 256$, which is unbearably slow – especially when compared to simple autoregressive calculation of log-likelihood that requires only a single forward pass.
Appendix E Pseudo log-likelihood estimation
@@
$\log p_{\bm{\theta}}(\bm{w})\approx\sum_{i=1}^{|{w}|}\log p_{\bm{\theta}}\bigl(w_{i}\mid\bm{c}$ $\oplus w_{0}\oplus\cdots\oplus w_{i-1}$ (9)
$\oplus\mathord{\texttt{mask}}$
@@
$\log p_{\bm{\theta}}(\bm{w})\approx\sum_{i=1}^{|\bm{w}|}\log p_{\bm{\theta}}\bigl(w_{i}\mid\bm{c}$ $\oplus w_{0}\oplus\cdots\oplus w_{i-1}$ (10)
$\oplus\texttt{mask}\oplus\cdots\oplus\texttt{mask}$
@@
where $n$ represents the number of $\operatorname{[MASK]}$ tokens. In our case, we take a combination of two different numbers of mask tokens (1 and 6), by taking the best score of the two for each task. The two values were chosen experimentally, more details on the results of each number of mask tokens can be found in Appendix H.
Appendix F Proof of left-shift closure
@@
The Restricted Access Sequence Processing language (RASP; Weiss et al., 2021) is a sequence processing language that uses two types of variables: sequence operators and selectors; and two types of operators: element-wise and select-aggregate operators. Valid programs in RASP are operations on sequence operators formed by a finite composition of element-wise and select-aggregate operators.
@@
For every valid program written in RASP, there exists an equivalent fully-bidirectional transformer model that computes the same per-position operation; see Weiss et al. (2021); Lindner et al. (2023).
@@
Theorem 1 implies that when we parameterize a masked-diffusion model with a shifted transformer, it is as expressive as the standard non-shifted parameterization. More specifically, masked diffusion is defined in Equation 4, and $p_{\bm{\theta}}(x_{i}\mid\bm{x}^{t})$ is typically implemented as a fully-bidirectional transformer model that outputs this probability at the $i$th position. When we set $\Sigma$ as our subword vocabulary, we get that the space of all possible transformer realizations of $p_{\bm{\theta}}(x_{i}\mid\bm{x}^{t})$ are the $\Sigma$-realizable functions $\mathcal{R}_{\Sigma}$ (Definition 2). Theorem 1 shows that if we instead expect the output at the (i1)(i-1)th position, we do not lose any expressivity. Thus, transformer-based dual-objective language models are a generalization of standard masked-diffusion language models. Note that the left-shift closure in Theorem 1 works up to the first token – which is guaranteed to be the special <s> token in Definition 2 as well as in the actual implementation.
Appendix G Validation loss curves
@@
Figure 6: Validation loss curves for 128 repetitions. These plots clearly demonstrate how training runs with high $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} (in red) overfit. Low $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} values are in blue.
@@
Figure 7: Validation loss curves for 4 repetitions. All losses monotonically decrease because overfitting is not a concern in this setting. High $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} values are plotted in red and low $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} values are shown in blue.
Appendix H Effects of number of mask tokens on the PLL
We first look at whether using a single number of mask tokens can lead to a good estimation of the PLL in general. For this, we evaluate five different models from 1 to 6 mask tokens and report the results in Tables 4, 4, 6, 6 and 7.
Table 3: PLL performance depending on the number of mask tokens. We show the PLL performance on the 9 tasks of the model trained with an equal $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} weight of masked-diffusion and AR and 32 repetitions with different number of masks. Best results per task are boldfaced.
@@ Appendix H Effects of number of mask tokens on the PLL
Table 4: PLL performance depending on the number of mask tokens. We show the PLL performance on the 9 tasks of the model trained with a 1 masked-diffusion to 7 autoregressive ratio (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=\nicefrac{{7}}{{8}}$) and 32 repetitions with different number of masks. Best results per task are boldfaced.
@@
Table 5: PLL performance depending on the number of mask tokens. We show the PLL performance on the 9 tasks of the model trained with a 7 masked-diffusion to 1 autoregressive ratio (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=\nicefrac{{1}}{{8}}$) and 32 repetitions with different number of masks. Best results per task are boldfaced.
@@
Table 6: PLL performance depending on the number of mask tokens. We show the PLL performance on the 9 tasks of the model trained with an equal ratio of masked-diffusion and AR (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=\nicefrac{{1}}{{2}}$) and 16 repetitions with different number of masks. Best results per task are boldfaced.
@@
Table 7: PLL performance depending on the number of mask tokens. We show the PLL performance on the 9 tasks of the model trained with an equal ratio of masked-diffusion and AR (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=\nicefrac{{1}}{{2}}$) and 64 repetitions with different number of masks. Best results per task are boldfaced.
@@
Table 8: PLL performance for combinations of one mask token and multi-mask token. Best results per model are boldfaced.
@@
Appendix I Total compute resources used for training
The training of all 50 language models used in this paper was conducted on the LUMI supercomputer, each language model was trained on 128 AMD MI250X GPUs (which is equivalent to 256 logical devices) using roughly 1 500 GPU hours. In total, the resources required for conducting all training runs equals to approximately 75 000 GPU hours.
Appendix J PLL versus masked diffusion
Table 9: Normalized PLL versus Masked-Diffusion evaluation. The scores for each task are normalized so that 0% corresponds to the random baseline and 100% is the perfect score. The best result for each task is in boldfaced. We evaluate a model trained with equal AR and masked-diffusion ratio (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=\nicefrac{{1}}{{2}}$) and 32 repetitions.
@@ Appendix I Total compute resources used for training
Appendix K Prefix-LM versus autoregressive-LM on optimal models.
@@
Table 10 shows that evaluating with the prefix mask almost always outperforms using the causal mask when the models are optimally trained. This is true in both the regular and constrained data settings.
Appendix L Detailed results of diffusion-masked evaluation
Table 11: The normalized PLL performance of selected models. We show the results on all nine evaluated tasks for three repetition values; each repetition group contains the results of the best-performing $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} and of the autoregressive-only model. The scores for each task are normalized so that 0% corresponds to random baseline and 100% is the perfect score. The best result for each dataset size is boldfaced.
@@
   Dual (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=\nicefrac{{3}}{{4}}$) 06.0 28.3 62.7 33.4 27.8 04.3 12.3 37.4 15.4 25.3
   Masked-Diffusion (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=0$)  -0.1 22.3 64.8 29.0 24.1 01.6 09.1 27.2 14.4 21.4
@@
   Dual (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=\nicefrac{{1}}{{8}}$) 02.8 23.3 63.5 30.5 25.0 02.1 12.8 31.8 15.2 23.0
@@
Table 11 shows similar trends to those found in Table 2. The notable exception being for BLiMP where the performances are similar between both models. Unlike the autoregressive models, the performance of the purely masked-diffusion models is similar to each other. This is partially due to the model not overfitting, but also to it not being sample efficient. On the other hand we see that for the dual-objective models, the performance significantly increases as we increase the training data set size.